Consider the case of Marston Manufacturing Company: Marston Manufacturing Company is considering a project that requires an iMarstons initial net investment outlay is $2,915,500 Suppose Marstons new equipment is expl pr $400,000 at the end of its fSuppose Marstons new equipment is expected to sell for $400,000 at the end of its four-year useful life, and at the same tim
Answer #1

After tax cost of Equipment = Purchase cost – Tax savings on Depreciation

= 3,570,000 – 3,570,000*25%

= $2,677,500

Initial Outlay = Cost of Equipment + Increase in current assets – Increase in current liabilities

= 2,677,500 + 680,000-272000

= $3,085,500

Termination Cash flows:

After tax salvage value = (400,000-0)*(1-25%) = $300,000

Recovery of Working Capital = $408,000

Termination Value = $708,000